OpenStudy (anonymous):

what is 5^x=372 solved for x

OpenStudy (anonymous):

\[b^x=A\\x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (solomonzelman):

log base 5 of both sides, that is first that comes to my mind

Directrix (directrix):